OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively.

OpenStudy (anonymous):

The nth term is : \[u_n=aq^n\] And we have : \[u_{2}=-12~~~u_5=768\] So we get : \[-12=aq^2~~~(1)\\768=aq^5~~~~(2)\] By dividing (2) over (1) we get : \[q^3=\frac{-768}{12}=-64\] So : \[q=\sqrt[3]{-64}=-4\] From (1) we have : \[12=aq^2\] So : \[a=\frac{q^2}{12}=\frac{(-4)^2}{12}=\frac{16}{12}=\frac43\] So the nth term is : \[u_n=\frac43\times(-4)^n\]

OpenStudy (anonymous):

Thank you so much!